### v1

@@ C.2 Benchmark calculations of $\mathcal{A}$ at real energies
In Eq. (99), the inhomogeneous term is always smooth in momentum space and spatially localized, but the kernel in the integration term is singular when $\mathrm{Im}\,(E)=0^{+}$, which requires dedicated treatments. Such treatment should apply to the equation with smooth sources, such as in computing on-shell $T$-matrix and the response-function-like calculations we set up by fixing the $E_{rel}$ parameter in the inhomogeneous term.
@@
To deal with singular kernel in Eq. (99), we follow the procedure of Ref. Matsuyama et al. (2007); Larson and Hetherington (1974) by carefully choosing the momentum mesh points for both $P$ and $P^{\prime}$ variable and by proper variable transformations around the singularities of the $\bar{Z}$ function. The other factor in the kernel, $\tau_{1}(E-P^{\prime 2}/(2\mu^{1}))$, has a simple pole located at $E-P^{\prime 2}/(2\mu^{1})\to-B_{2}$. This pole can be dealt with by a simple pole subtraction method commonly used in solving two-body Lippmann-Schwinger equation (see, e.g., Sec. 18.3 in Ref. Landau (1996)).
@@
Specifically, in the integral equation, treating $f(P^{\prime};P,E)\equiv\bar{Z}(P,P^{\prime},E,E,1,1)$ as a function of $P^{\prime}$ variable with a fixed $P$ and $E$, requires careful analysis. For $E<0$, the $f(P^{\prime};P,E)$ is smooth. However, when $E>0$, $f(P^{\prime};P,E)$ becomes singular near two points (collectively labeled as $P_{ls}\equiv|\sqrt{ME-3P^{2}/4}\pm P/2|$): $\mathrm{Re}\,f$ diverges logarithmically and $\mathrm{Im}\,f$ becomes discontinuous. Therefore, we first decompose the mesh for $P^{\prime}$ (and $P$) into several intervals Matsuyama et al. (2007); Larson and Hetherington (1974): $[0,P_{1}]$, $[P_{1},P_{2}]$, $[P_{2},P_{3}]$, $[P_{3},P_{rel}]$, and $[P_{rel},P_{max}]$ with $P_{1}\equiv\sqrt{M_{N}E/3}$, $P_{2}\equiv\sqrt{M_{N}E}$, $P_{3}\equiv\sqrt{4M_{N}E/3}$, $P_{rel}=\sqrt{4M_{N}(E+B_{2})/3}$, and $P_{max}$ as the UV cut off. When there is no bound dimer, $B_{2}=0$ is assumed in this setup of the intervals. Note that these two singularities at $P_{ls}$ are always in two different intervals. In the intervals without the two singularities, we apply the Lagrange function-based interpolant, which assigns momentum mesh in each interval according to the Gaussian quadrature Baye (2015). To deal with the intervals with the two singularities, we make a variable transformation with $P^{\prime}-P_{ls}=t^{3}$ around the singularities, with which the kernel becomes smooth in terms of $t$. Then, the Gaussian quadrature mesh in terms of the $t$ variable informs the momentum mesh.
@@
Moreover, as the solution to the Eq. (99), $F(P)$ has a branch point at $P_{3}$, when $E\geq 0$. Near that point, $F(P)\sim\sqrt{P_{3}-P}$ Matsuyama et al. (2007); Larson and Hetherington (1974). We apply another variable transformation Matsuyama et al. (2007) so that when $P<P_{3}$, $t\sim\sqrt{P_{3}-P}$ and when $P>P_{3}$, $t\sim\sqrt{P-P_{3}}$ with proper normalizations. Again, the Gaussian quadrature mesh for $t$ provides the corresponding momentum mesh.
@@
It is worth pointing out that in Refs. Matsuyama et al. (2007); Larson and Hetherington (1974), there were no consideration of three-body interactions. However, adding this interaction does not add new singularities in $f(P^{\prime})$ and no new branch points in the solution $F(P)$ to the integral equation.
@@ C.3 Benchmark calculations for bound and resonance states
Here, we discuss the benchmark calculations of the locations of physical states. In order to compute bound state eigenenergies, we look for the zeros of the kernel, $\bar{Z}(P,P^{\prime},E,E,1,1)$, as a function of $E$ with $E$ restricted to be real and below the particle-dimer threshold. The kernel, numerically, is a matrix in two-dimensional momentum mesh for $P$ and $P^{\prime}$. A simple Gaussian quadrature mesh suffices since the kernel varies smoothly in the bound state region when changing $P$ and $P^{\prime}$. We then look for the zeros of the matrix’s determinant.
To compute the resonance location, we could have followed the established method using contour deformation or complex-scaling method. However, for simplicity, we apply the rational approximation as an analytical continuation method to continue $\mathcal{A}$ from the benchmark calculation results on the real energy axis ($\mathrm{Im}\,E\to 0^{+}$) to the region below the real axis in the complex $E$ plane. The rational approximation is
@@
In this expression, the first term is the resonance pole contribution—assuming we know the existence of single resonance, the second is the pole contribution from possible three-body bound states, and the last term summing up the BCT poles. For a given continuous BCT that starts from the corresponding branching point $E=E_{th}$ and has a relative angle ($\theta_{\mathrm{BCT}}$) from the positive real axis, the discretized BCT poles are distributed on that line according to the tapered exponential distribution Trefethen et al. (2021). The location distribution is parameterized with $E_{th}$, $c_{0}$, $c_{1}$, $N_{\mathrm{BCT}}$, and $\theta_{\mathrm{BCT}}$:
@@
Then, we approximate $\mathcal{A}(E)\approx f(E)$ and look for the best fits that minimize the difference between the two at chosen $E^{tr}_{\alpha}$. It should be emphasized that here $E^{tr}_{\alpha}$ are on the upper rim of the real energy axis ($\mathrm{Im}\,E^{tr}_{\alpha}\to 0^{+}$) and spread out over large energy intervals.
The minimization is performed in two steps. For a given $E_{r}$, the residues $w_{i}$ and $w_{r}$ can be solved by applying the pseudo-inverse method to solve the over-constrained linear equation system: $f(E^{tr}_{\alpha})=\mathcal{A}(E^{tr}_{\alpha})$. Then, we compute the loss function defined as $\sum_{\alpha}|f(E^{tr}_{\alpha})-\mathcal{A}(E^{tr}_{\alpha})|$. The best $E_{r}$ is identified if it minimizes the loss function. The robustness of the minimization is checked by varying the parameters for controlling the distribution of the BCT poles and looking for the consistency between these different fits.

### v2

@@ C.2 Benchmark calculations of $\mathcal{A}$ at real energies
In Eq. (101), the inhomogeneous term is always smooth in momentum space and spatially localized, but the kernel in the integration term is singular when $\mathrm{Im}\,(E)=0^{+}$, which requires dedicated treatments. Such treatment should apply to the equation with smooth sources, such as in computing on-shell $T$-matrix and the response-function-like calculations I set up by fixing the $E_{\mathrm{rel}}$ parameter in the inhomogeneous term.
@@
To deal with singular kernel in Eq. (101), I follow the procedure of Ref. [118, 119] by carefully choosing the momentum mesh points for both $P$ and $P^{\prime}$ variable and by proper variable transformations around the singularities of the $\bar{Z}$ function. The other factor in the kernel, $\tau_{1}(E-P^{\prime 2}/(2\mu^{1}))$, has a simple pole located at $E-P^{\prime 2}/(2\mu^{1})\to-B_{2}$. This pole can be dealt with by a simple pole subtraction method commonly used in solving two-body Lippmann-Schwinger equation (see, e.g., Sec. 18.3 in Ref. [120]).
@@
Specifically, in the integral equation, treating $f(P^{\prime};P,E)\equiv\bar{Z}(P,P^{\prime},E,E,1,1)$ as a function of $P^{\prime}$ variable with a fixed $P$ and $E$, requires careful analysis. For $E<0$, the $f(P^{\prime};P,E)$ is smooth. However, when $E>0$, $f(P^{\prime};P,E)$ becomes singular near two points (collectively labeled as $P_{ls}\equiv|\sqrt{ME-3P^{2}/4}\pm P/2|$): $\mathrm{Re}\,f$ diverges logarithmically and $\mathrm{Im}\,f$ becomes discontinuous. Therefore, I first decompose the mesh for $P^{\prime}$ (and $P$) into several intervals [118, 119]: $0,P_{1}]$, $P_{1},P_{2}]$, $P_{2},P_{3}]$, $P_{3},P_{\mathrm{rel}}]$, and $P_{\mathrm{rel}},P_{max}]$ with $P_{1}\equiv\sqrt{M_{N}E/3}$, $P_{2}\equiv\sqrt{M_{N}E}$, $P_{3}\equiv\sqrt{4M_{N}E/3}$, $P_{\mathrm{rel}}=\sqrt{4M_{N}(E+B_{2})/3}$, and $P_{max}$ as the UV cut off. When there is no bound dimer, $B_{2}=0$ is assumed in this setup of the intervals. Note that these two singularities at $P_{ls}$ are always in two different intervals. In the intervals without the two singularities, I apply the Lagrange function-based interpolant, which assigns momentum mesh in each interval according to the Gaussian quadrature [80]. To deal with the intervals with the two singularities, I make a variable transformation with $P^{\prime}-P_{ls}=t^{3}$ around the singularities, with which the kernel becomes smooth in terms of $t$. Then, the Gaussian quadrature mesh in terms of the $t$ variable informs the momentum mesh.
@@
Moreover, as the solution to the Eq. (101), $F(P)$ has a branch point at $P_{3}$, when $E\geq 0$. Near that point, $F(P)\sim\sqrt{P_{3}-P}$ [118, 119]. I apply another variable transformation [118] so that when $P<P_{3}$, $t\sim\sqrt{P_{3}-P}$ and when $P>P_{3}$, $t\sim\sqrt{P-P_{3}}$ with proper normalizations. Again, the Gaussian quadrature mesh for $t$ provides the corresponding momentum mesh.
@@
It is worth pointing out that in Refs. [118, 119], there were no consideration of three-body interactions. However, adding this interaction does not add new singularities in $f(P^{\prime})$ and no new branch points in the solution $F(P)$ to the integral equation.
@@ C.3 Benchmark calculations for bound and resonance states
Here, I discuss the benchmark calculations of the locations of physical states. In order to compute bound state eigenenergies, I look for the zeros of the kernel, $\bar{Z}(P,P^{\prime},E,E,1,1)$, as a function of $E$ with $E$ restricted to be real and below the particle-dimer threshold. The kernel, numerically, is a matrix in two-dimensional momentum mesh for $P$ and $P^{\prime}$. A simple Gaussian quadrature mesh suffices since the kernel varies smoothly in the bound state region when changing $P$ and $P^{\prime}$. I then look for the zeros of the matrix’s determinant.
To compute the resonance location, I could have followed the established method using contour deformation or complex-scaling method. However, for simplicity, I apply the rational approximation as an analytical continuation method to continue $\mathcal{A}$ from the benchmark calculation results on the real energy axis ($\mathrm{Im}\,E\to 0^{+}$) to the region below the real axis in the complex $E$ plane. The rational approximation is
@@
In this expression, the first term is the resonance pole contribution—assuming the existence of a single resonance, the second is the pole contribution from possible three-body bound states, and the last term summing up the BCT poles. For a given continuous BCT that starts from the corresponding branching point $E=E_{th}$ and has a relative angle ($\theta_{\mathrm{BCT}}$) from the positive real axis, the discretized BCT poles are distributed on that line according to the tapered exponential distribution [96]. The location distribution is parameterized with $E_{th}$, $c_{0}$, $c_{1}$, $N_{\mathrm{BCT}}$, and $\theta_{\mathrm{BCT}}$:
@@
Then, I approximate $\mathcal{A}(E)\approx f(E)$ and look for the best fits that minimize the difference between the two at chosen $E^{\mathrm{tr}}_{\alpha}$. It should be emphasized that here $E^{\mathrm{tr}}_{\alpha}$ are on the upper rim of the real energy axis ($\mathrm{Im}\,E^{\mathrm{tr}}_{\alpha}\to 0^{+}$) and spread out over large energy intervals.
The minimization is performed in two steps. For a given $E_{r}$, the residues $w_{i}$ and $w_{r}$ can be solved by applying the pseudo-inverse method to solve the over-constrained linear equation system: $f(E^{\mathrm{tr}}_{\alpha})=\mathcal{A}(E^{\mathrm{tr}}_{\alpha})$. Then, I compute the loss function defined as $\sum_{\alpha}|f(E^{\mathrm{tr}}_{\alpha})-\mathcal{A}(E^{\mathrm{tr}}_{\alpha})|$. The best $E_{r}$ is identified if it minimizes the loss function. The robustness of the minimization is checked by varying the parameters for controlling the distribution of the BCT poles and looking for the consistency between these different fits.